# Hierarchical Memory for Long Video QA

### v1

@@ 1 Introduction
Most existing large video-language models face challenges when performing long video question-answering upon user queries [3, 2, 6, 10]. The main reason is that: visual tokens between consecutive frames are heavy and redundant without effective compression, making it impossible to save all visual features in limited GPU Memory (VRAM), as well as significantly increasing the decoding latency of language model. To better address this issue, we adopt a hierarchical memory mechanism named STAR Memory, which is proposed recently in Flash-VStream [11]. By further fine-tuning the pretrained weight released by Flash-VStream on the training branch of MovieChat-1K, we successfully leverage the general video compression and understanding ability of STAR Memory, as well as the domain-specific knowledge of MovieChat-1K. The strong performance of our model demonstrates the effectiveness of the hierarchical memory method, namely the STAR Memory, in long video question-answering tasks.

### v2

@@ 1 Introduction
Recently, natural language is becoming a general interface for various tasks and modalities [1, 5, 10, 23, 17, 18] Most existing large video-language models face challenges when performing long video question-answering upon user queries [6, 4, 12, 16]. The main reason is that: visual tokens between consecutive frames are heavy and redundant without effective compression, making it impossible to save all visual features in limited GPU Memory (VRAM), as well as significantly increasing the decoding latency of language model. Some previous works solve this by feature compression or pruning [21, 20, 19, 9], and achieve real-time performance [22, 3, 11]. To better address this issue, we adopt a hierarchical memory mechanism named STAR Memory, which is proposed recently in Flash-VStream [22]. By further fine-tuning the pretrained weight released by Flash-VStream on the training branch of MovieChat-1K, we successfully leverage the general video compression and understanding ability of STAR Memory, as well as the domain-specific knowledge of MovieChat-1K. The strong performance of our model demonstrates the effectiveness of the hierarchical memory method, namely the STAR Memory, in long video question-answering tasks.